# Affordance-based Robot Manipulation with Flow Matching

### v1

@@ Abstract
We present a framework for assistive robot manipulation, which focuses on two fundamental challenges: first, efficiently adapting large-scale models to downstream scene affordance understanding tasks, especially in daily living scenarios where gathering multi-task data involving humans requires strenuous effort; second, effectively learning robot trajectories by grounding the visual affordance model. We tackle the first challenge by employing a parameter-efficient prompt tuning method that prepends learnable text prompts to the frozen vision model to predict manipulation affordances in multi-task scenarios. Then we propose to learn robot trajectories guided by affordances in a supervised Flow Matching method. Flow matching represents a robot visuomotor policy as a conditional process of flowing random waypoints to desired robot trajectories. Finally, we introduce a real-world dataset with 10 tasks across Activities of Daily Living to test our framework. Our extensive evaluation highlights that the proposed prompt tuning method for learning manipulation affordance with language prompter achieves competitive performance and even outperforms other finetuning protocols across data scales, while satisfying parameter efficiency. Learning multi-task robot trajectories with a single flow matching policy also leads to consistently better performance than alternative behavior cloning methods, especially given multimodal robot action distributions. Our framework seamlessly unifies affordance model learning and trajectory generation with flow matching for robot manipulation. https://hri-eu.github.io/flow-matching-policy/.
@@ 1 Introduction
Recent advances in vision-language models (VLMs) present unprecedented opportunities to solve robot manipulation problems. Attempts in the field have focused on three primary aspects: 1) End-to-end learning manipulation from scratch. These approaches (Padalkar et al., 2023) make the least assumptions on tasks and are formulated in language-image-to-action prediction models. 2) Off-the-shelf-vision-language models for robot manipulation. Such line of works have explored using pre-trained VLMs with prompt engineering in various contexts of robot motion learning, including reward design for reinforcement learning (Ma et al., 2023), python coding (Liang et al., 2023), joint actions (Wang et al., 2023), etc. 3) Intermediate substrate to bridge high-level language-image instructions and low-level robot policies. These works usually introduce some form of prior derived from human knowledge as an intermediate stage, including affordances (Huang et al., 2023), pre-trained visual representations (Xiao et al., 2022), primitive skills (Ingelhag et al., 2024), etc. In this paper, we follow the third line of work. We propose to learn manipulation affordances, leveraging finetuning vision-language foundation models with a wealth of internalized knowledge. Then the affordance model is used for learning low-level robot policies, and thus helps to alleviate the sample inefficiency problem of end-to-end learning from scratch.
@@
Extracting affordance knowledge has long inspired the robot community (Xiao et al., 2022). We particularly focus on affordance learning in multi-task scenarios with text prompting. As shown in Fig. 2, given the same visual scene but with different language instructions, we aim to extract different affordances for robot policy learning through our proposed model. To leverage the ability of pre-trained foundation models while avoiding expensive computational constraints, recent works have explored parameter-efficiently fine-tuning (PEFT) large vision-language models for various visual recognition applications (Jia et al., 2022). PEFT could be mainly categorized into two main groups: adapter-based methods (e. g., LoRA (Hu et al., 2021)) and prompt-based methods (Liu et al., 2023). One representative line of such research has concentrated on prompt tuning methods, which prepend learnable prompts to the input of a large frozen pre-trained model and optimize them via gradients during finetuning. Studies on randomly-generalized trainable prompts (Jia et al., 2022) for universal use or condition-admitted prompt variables (Sohn et al., 2023) for better specific task performance have been both explored. It has shown that prompt tuning could match the performance of full finetuning but with substantially less parameter storage in various domains, including visual tracking (Zhu et al., 2023) and cross-domain tasks such as language-dance assessment (Zhong and Demiris, 2024). Inspired in part by the notion of human cognitive penetrability mechanism (Maier and Abdel Rahman, 2019) that uses linguistic knowledge to tune ongoing visual processing, we incorporate learnable text-conditioned prompts in a frozen vision foundation model to learn instruction-relevant manipulation affordance.
The subsequent challenge involves deploying the visual affordance across various robot manipulation learning paradigms. From the traditional behavior cloning with convolutional networks (Zhang and Demiris, 2022) to transformer-based learning structures (Shridhar et al., 2023), extensive research has modeled robot action trajectories from visual scenes. A recent line of works builds on successes in diffusion models (Chi et al., 2023) to generate motion trajectories, to capture multimodal action distributions and stable training behavior. Flow Matching is another novel generative method. Sharing theoretical similarities with stochastic Denoising Diffusion Probabilistic Models, flow matching aims to regress onto a deterministic vector field to flow samples toward the target distribution. It has proven that the simplicity of flow matching objectives allows favorable performance in stable training and generation quality compared to solving complex stochastic differential equations in diffusion models. We extend flow matching to the robotics domain. This is the first attempt to leverage flow matching in real-world robot manipulation from visual affordance. As shown in Fig. 1 and Fig. 2, the proposed method would flow the random waypoints to the desired trajectories based on various affordances in a single flow matching policy.
We also construct a real-world dataset with 10 tasks across Activities of Daily Living to test the proposed method. The novelty of our dataset is that it contains same scenarios but with multi-task affordance and robot trajectories. Experimental evaluation empirically demonstrates that the prompt tuning method for learning affordances achieves performance competitive, and sometimes beyond other finetuning protocols across data scales, vision-language fusion architectures, and prompt variants. Furthermore, we showcase that flow matching policy attains favorable performance in stable training and generation quality amongst competing methods of robot behavior cloning. Note that our goal is not to achieve the state-of-the-art general robot manipulation performance, but instead to broadly explore a new paradigm of efficiently adapting VLMs for affordance learning, and robot policy for multimodal action distributions.
@@
    Empirical results investigating the advantages of the proposed algorithm on our constructed real-world robot multi-task dataset.
@@ 2 Related Work
2.0.1 Robot Learning from Demonstration
Imitation learning has been a common paradigm for robots which requires simulated or real-world demonstration data collection (Zhang and Demiris, 2023). To improve data efficiency, extensive works have been proposed to learn robot policies on the top of visual representations (Liu et al., 2024) instead of end-to-end raw images (Goyal et al., 2023). Keypoints or affordance heatmaps are often used to provide contact information as visual representation (Liu et al., 2024). The concept of affordance has been introduced in (Gibson, 2014), referring to the ability to perform certain actions with objects in the context of a given scene. This paper concentrates on using affordances to guide the low level robot manipulation. In terms of network architectures for robot learning, prior works have successfully investigated convolutional networks (Zhang and Demiris, 2022), Transformers (Shridhar et al., 2023), generative adversarial networks (Ho and Ermon, 2016), Energy-Based Models (Florence et al., 2022), etc. However, the collected data is usually expected to be non-convex and multi-modal due to the variability in human demonstrations. Recent works have addressed this problem by reformulating the robot policy as a generative process. Diffusion policy (Chi et al., 2023) has emerged as a powerful class of generative models for behavior cloning by representing a robot’s visuomotor policy as a conditional denoising diffusion process. In this work, we investigate Flow Matching (Lipman et al., 2022), a novel generative model that has demonstrated its superiority in image generation, but is much less explored in robotics domains.
2.0.2 Parameter-Efficient Finetuning
@@
2.0.3 Flow Matching in Robotics
@@ 3.1.1 Shallow Architecture
The vision transformer layer takes the image patch embeddings $\bm{E}_{0}$ as input, and passes through various layers $\bm{L}_{i}^{v}$ to achieve vision features $\bm{E}_{i}$, where $\bm{E}_{i}\in\mathbb{R}^{M\times C}$ and $C$ is the channel dimension.
@@ 3.2 Flow Matching Policy
In this session, we build the robot behavioral cloning policy as a generative process of Flow Matching, which constructs a flow vector that continuously transforms a source probability distribution towards a destination distribution. Flow Matching leverages an ordinary differential equation to deterministically mold data distribution, contrasting with diffusion policy which is based on a stochastic differential equation through introducing noise.
Figure 4: Framework of Flow Matching Policy. (a) At each time step, the policy predicts flow vectors for each waypoint conditioned on visual affordance observation data as well as the waypoints at the current timestep. (b-c) Flow matching transforms random waypoints (green) to the target trajectory (purple) from timestep 0 to 1. The lines in (b) denote the flow vectors.
@@ 3.2.1 Flow Matching Model
Given a probability density path $p_{t}(\bm{x})$ and a corresponding vector field $\bm{u}_{t}(\bm{x})$, the objective loss of flow matching could be described as:
$\mathcal{L}_{\text{FM}}(\bm{\theta})=\mathbb{E}_{t,p_{t}(\bm{x})}\left\|\mathbf{v}_{t}(\bm{x},\bm{\theta})-\mathbf{u}_{t}(\bm{x})\right\|^{2}$ (1)
where $\bm{x}\sim p_{t}(\bm{x})$, $t\sim\mathcal{U}[0,1]$ (uniform distribution). The aim of flow matching is to regress $\mathbf{u}_{t}(\bm{x})$ with a time-dependent vector field of flow $\mathbf{v}_{t}(\bm{x},\bm{\theta})$ parameterized as a neural network with weights $\bm{\theta}$. $\mathbf{u}_{t}(\bm{x})$ can be represented as
$\mathbf{u}_{t}(\bm{x})=\bm{x}_{1}-\bm{x}_{0}\qquad\bm{x}_{0}\sim p_{0},\bm{x}_{1}\sim p_{1}$
$p_{0}$ represents a simple base density at time $t=0$, $p_{1}$ denotes the target complicated distribution at time $t=1$, $\bm{x}_{0}$ and $\bm{x}_{1}$ are the corresponding samplings. $\mathbf{v}_{t}(\bm{x},\bm{\theta})$ could be described as:
$\mathbf{v}_{t}(\bm{x},\bm{\theta})=f_{\bm{\theta}}(\bm{x}_{t},t)$ (2)
We define $\bm{x}_{t}$ as the linear interpolation between $\bm{x}_{0}$ and $\bm{x}_{1}$ with respect to time $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, following the Optimal Transport theory (Peyré et al., 2019). And $f_{\bm{\theta}}$ is a network of the flow model. Thus Equation (1) could be reformatted as:
$\mathcal{L}_{\text{FM}}(\bm{\theta})=\mathbb{E}_{t,\sim p_{0},\sim p_{1}}\left\|f_{\bm{\theta}}(\bm{x}_{t},t)-(\bm{x}_{1}-\bm{x}_{0}))\right\|^{2}$ (3)
@@
Algorithm 1 Flow Matching Policy
affordance observation $\bm{O}_{t}$, demonstrated robot trajectory $\bm{x}_{1}$, source random waypoints $p_{0}$
flow $\mathbf{v}_{t}(\bm{x}|\bm{o}_{t})$

### v2

@@ Abstract
We present a framework for assistive robot manipulation, which focuses on two fundamental challenges: first, efficiently adapting large-scale models to downstream scene affordance understanding tasks, especially in daily living scenarios where gathering multi-task data involving humans requires strenuous effort; second, effectively learning robot trajectories by grounding the visual affordance model. We tackle the first challenge by employing a parameter-efficient prompt tuning method that prepends learnable text prompts to the frozen vision model to predict manipulation affordances in multi-task scenarios. Then we propose to learn robot trajectories guided by affordances in a supervised Flow Matching method. Flow matching represents a robot visuomotor policy as a conditional process of flowing random waypoints to desired robot trajectories. Finally, we introduce a real-world dataset with 10 tasks across Activities of Daily Living to test our framework. Our extensive evaluation highlights that the proposed prompt tuning method for learning manipulation affordance with language prompter achieves competitive performance and even outperforms other finetuning protocols across data scales, while satisfying parameter efficiency. Learning multi-task robot trajectories with flow matching policy also leads to consistently better generalization performance and faster inference than alternative behavior cloning methods, especially given multimodal robot action distributions. Our framework seamlessly unifies affordance model learning and trajectory generation with flow matching for robot manipulation. https://hri-eu.github.io/flow-matching-policy/.
@@ 1 Introduction
Recent advances in vision-language models (VLMs) present unprecedented opportunities to solve robot manipulation problems. Attempts in the field have focused on three primary aspects: 1) End-to-end learning manipulation from scratch. These approaches (Padalkar et al., 2023) make the least assumptions on tasks and are formulated in language-image-to-action prediction models. 2) Off-the-shelf-vision-language models for robot manipulation. Such line of works have explored using pre-trained VLMs with prompt engineering in various contexts of robot motion learning, including reward design for reinforcement learning (Ma et al., 2023), python coding (Liang et al., 2023), joint actions (Wang et al., 2023), etc. 3) Intermediate substrate to bridge high-level language-image instructions and low-level robot policies. These works usually introduce some form of prior derived from human knowledge as an intermediate stage, including affordances (Huang et al., 2023), pre-trained visual representations (Xiao et al., 2022), primitive skills (Ingelhag et al., 2024), etc. In this paper, we follow the third line of work to unify an affordance model and low-level robot policies. which helps to alleviate the sample inefficiency problem of end-to-end learning from scratch.
Extracting affordance knowledge has long inspired the robot community (Xiao et al., 2022). Humans heavily rely on visual affordances to efficiently perform day-to-day tasks across environments. The concept of affordance has been introduced in (Gibson, 2014), referring to the ability to perform certain actions with objects in the context of a given scene. In this work, affordance focuses on the areas of an object that permit or enable specific actions. Affordance models allow more knowledge transfer across different objects with similar affordances, and more intuitive interaction with environments, compared to other intermediate substrates, including keypoints or rewards. We particularly focus on affordance learning in multi-task scenarios with text prompting. As shown in Fig. 1, given the same visual scene but with different language instructions, we aim to extract different affordances for robot policy learning through our proposed model. To leverage the ability of pre-trained foundation models while avoiding expensive computational constraints, recent works have explored parameter-efficiently fine-tuning (PEFT) large vision-language models for various visual recognition applications (Jia et al., 2022). PEFT could be mainly categorized into two main groups: adapter-based methods (e. g., LoRA (Hu et al., 2021)) and prompt-based methods (Liu et al., 2023). One representative line of such research has concentrated on prompt tuning methods, which prepend learnable prompts to the input of a large frozen pre-trained model and optimize them via gradients during finetuning. Studies on randomly-generalized trainable prompts (Jia et al., 2022) for universal use or condition-admitted prompt variables (Sohn et al., 2023) for better specific task performance have been both explored. It has shown that prompt tuning could match the performance of full finetuning but with substantially less parameter storage in various domains, including visual tracking (Zhu et al., 2023) and cross-domain tasks such as language-dance assessment (Zhong and Demiris, 2024). Inspired in part by the notion of human cognitive penetrability mechanism (Maier and Abdel Rahman, 2019) that uses linguistic knowledge to tune ongoing visual processing, we target to incorporate learnable text-conditioned prompts into any vision foundation model while keeping it frozen, preserving its visual understanding capabilities, to learn instruction-relevant manipulation affordance.
The subsequent challenge involves deploying the visual affordance across various robot manipulation learning paradigms. From the traditional behavior cloning with convolutional networks (Zhang and Demiris, 2022) to transformer-based learning structures (Shridhar et al., 2023), extensive research has modeled robot action trajectories from visual scenes. A recent line of works builds on successes in diffusion models (Chi et al., 2023) to generate motion trajectories to capture multimodal action distributions. Flow Matching is another novel generative method. Sharing theoretical similarities with stochastic Denoising Diffusion Probabilistic Models (DDPM), flow matching aims to regress onto a deterministic vector field to flow samples toward the target distribution. It has proven that the simplicity of flow matching objectives allows favorable performance in stable training and generation quality compared to solving complex stochastic differential equations in DDPM. We extend flow matching to the robotics domain. As shown in Fig. 1 and Fig. 1, the proposed method would flow the random waypoints to the desired trajectories based on various affordances in a single flow matching policy.
@@
We also construct a real-world dataset with 10 tasks across Activities of Daily Living to test the proposed method. The novelty of our dataset is that it contains the same scenarios but with multi-task affordance and robot trajectories. Experimental evaluation on our dataset and other benchmarks empirically demonstrates that the prompt tuning method for learning affordances achieves performance competitive, and sometimes beyond other finetuning protocols across data scales, vision-language fusion architectures, and prompt variants. Furthermore, we showcase that flow matching policy attains favorable performance in stable training, generation quality and computational efficiency amongst competing methods of robot behavior cloning.
This is the first attempt to ground VLM-based affordance with flow matching for real-world robot manipulation. Note that our goal is not to achieve the state-of-the-art general robot manipulation performance, but instead to broadly explore a new paradigm of efficiently adapting VLMs for affordance learning, and robot policy for multimodal action distributions.
@@
    Empirical results show that flow matching policy leads to consistently better generalization performance and faster inference than alternative behavior cloning methods, including diffusion policy.
@@ 2 Related Work
2.1 Robot Learning from Demonstration
Imitation learning has been a common paradigm for robots which requires simulated or real-world demonstration data collection (Zhang and Demiris, 2023). To improve data efficiency, extensive works have been proposed to learn robot policies on the top of visual representations (Liu et al., 2024) instead of end-to-end raw images (Goyal et al., 2023). Keypoints or affordance heatmaps are often used to provide contact information as visual representation (Liu et al., 2024). This paper concentrates on using affordances to guide the low-level robot manipulation. In terms of network architectures for robot learning, prior works have successfully investigated convolutional networks (Zhang and Demiris, 2022), Transformers (Shridhar et al., 2023), generative adversarial networks (Ho and Ermon, 2016), Energy-Based Models (Florence et al., 2022), etc. However, the collected data is usually expected to be non-convex and multi-modal due to the variability in human demonstrations. Recent works have addressed this problem by reformulating the robot policy as a generative process. Diffusion policy (Chi et al., 2023) has emerged as a powerful class of generative models for behavior cloning by representing a robot’s visuomotor policy as a conditional denoising diffusion process. In this work, we investigate Flow Matching (Lipman et al., 2022), a novel generative model that has demonstrated its superiority in image generation, but is much less explored in robotics domains.
2.2 Parameter-Efficient Finetuning
@@
2.3 Flow Matching in Robotics
@@ 3.1.1 Shallow Architecture
The vision transformer layer takes the image patch embeddings $\bm{E}_{0}$ as input and passes through various layers $\bm{L}_{i}^{v}$ to achieve vision features $\bm{E}_{i}$, where $\bm{E}_{i}\in\mathbb{R}^{M\times C}$ and $C$ is the channel dimension.
@@ 3.1.3 Implementation Details
Our goal is to integrate textual representations into any vision encoder while keeping it frozen, preserving its visual understanding capabilities. Thus, we have chosen the most basic vision backbone, a pretrained ViT-B-16 transformer. The text layer adopts the classic CLIP setup, including Multiheaded Self-Attention, Feed-Forward Networks with LayerNorm and residual connections. As suggested by MAE (He et al., 2022), the decoder is only used for downstream tasks and could be flexible and lightweight. Thus we use one single transformer decoder layer.
@@ 3.2 Flow Matching Policy
We build the robot behavioral cloning policy as a generative process of Flow Matching, which constructs a flow vector that continuously transforms a source probability distribution toward a destination distribution. Flow Matching leverages an ordinary differential equation to deterministically mold data distribution, contrasting with diffusion policy which is based on a stochastic differential equation through introducing noise.
@@ 3.2.1 Flow Matching Model
Given a conditional probability density path $p_{t}(\bm{x}|\bm{z})$ and a corresponding conditional vector field $\bm{u}_{t}(\bm{x}|\bm{z})$, the objective loss of flow matching could be described as:
$\mathcal{L}_{\text{FM}}(\bm{\theta})=\mathbb{E}_{t,q(\bm{z}),p_{t}(\bm{x}|\bm{z})}\left\|\mathbf{v}_{t}(\bm{x},\bm{\theta})-\mathbf{u}_{t}(\bm{x}|\bm{z})\right\|^{2}$ (1)
where $\bm{x}\sim p_{t}(\bm{x}|\bm{z})$, $t\sim\mathcal{U}[0,1]$ (uniform distribution). Flow matching aims to regress $\mathbf{u}_{t}(\bm{x}|\bm{z})$ with a time-dependent vector field of flow $\mathbf{v}_{t}(\bm{x},\bm{\theta})$ parameterized as a neural network with weights $\bm{\theta}$. $\mathbf{u}_{t}(\bm{x}|\bm{z})$ can be further simplified as:
$\mathbf{u}_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}\qquad\bm{x}_{0}\sim p_{0},\bm{x}_{1}\sim p_{1}$
$p_{0}$ represents a simple base density at time $t=0$, $p_{1}$ denotes the target complicated distribution at time $t=1$, $\bm{x}_{0}$ and $\bm{x}_{1}$ are the corresponding samplings. $\mathbf{v}_{t}(\bm{x},\bm{\theta})$ could be described as
$\mathbf{v}_{t}(\bm{x},\bm{\theta})=v_{\bm{\theta}}(\bm{x}_{t},t)$ (2)
We define $\bm{x}_{t}$ as the linear interpolation between $\bm{x}_{0}$ and $\bm{x}_{1}$ with respect to time $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, following the Optimal Transport theory (Peyré et al., 2019). And $v_{\bm{\theta}}$ is a network of the flow model. Thus Equation (1) could be reformatted as
$\mathcal{L}_{\text{FM}}(\bm{\theta})=\mathbb{E}_{t,\sim p_{0},\sim p_{1}}\left\|v_{\bm{\theta}}(\bm{x}_{t},t)-(\bm{x}_{1}-\bm{x}_{0}))\right\|^{2}$ (3)
@@
Figure 4: Framework of Flow Matching Policy. (a) At each time step, the policy predicts flow vectors for each waypoint conditioned on visual observation data as well as the waypoints at the current timestep. (b-c) Visualization of the inference phase using Flow matching to transform random waypoints (green) to the target action trajectory (purple) from timestep 0 to 1. The lines in (b) denote the flow vectors.
3.2.2 Flow Matching for Visuomotor Policy Learning
We extend flow matching to learn robot visuomotor policies. This requires two modifications in the formulation: i) modeling the flow estimation conditioned on input observations $\bm{o}$; ii) changing the output $\bm{x}$ to represent robot actions. Fig. 4 illustrates our model structures.
Visual observation Conditioning: We modify Equation (2) to allow the model to predict actions conditioned on observations:
$\mathbf{v}_{t}(\bm{x}|\bm{o})=v_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})$
The visual embeddings $\bm{o}$ are obtained through ResNet (He et al., 2016). We also evaluate the Transformer structure in our experiments. Various types of inputs include state-based inputs, RGB images, and visual affordances.
Closed-loop action trajectory prediction: We execute the action trajectory prediction obtained by our flow matching model for a fixed duration before replanning. At each step, the policy takes the observation data $\bm{o}$ as input and predicts $Tp$ steps of actions, of which $Ta$ steps of actions are executed on the robot without re-planning. $Tp$ is the action prediction horizon and $Ta$ is the action execution horizon. The policy predicts flow vectors $\mathbf{v}_{t}$ conditioned on visual observation data $\bm{o}$ with Feature-wise Linear Modulation (FiLM) (Perez et al., 2018) as well as the interpolated waypoints $\bm{x}_{t}$. The flow model $f_{\bm{\theta}}$ is represented with U-Net (Ronneberger et al., 2015). The whole training process of the flow matching policy is illustrated in Algorithm 1. In the subsequent experiments, we will further study multiple model variants, including transformer-based structure, trajectory horizon, etc.
In our case of robot manipulation, $\bm{x}_{1}$ in Equation (3) represents the demonstration robot action trajectories. $\bm{x}_{0}$ is the random generated waypoints following a multivariate normal distribution $\bm{x}_{0}\sim\mathcal{N}(0,I)$. $\bm{x}$ here could denote 6D robot end-effector trajectories or robot joint actions.
Inference: For the inference procedure, random waypoints are sampled from the source distribution and then flowed into the target trajectory by estimating the flow from $t=0$ to $t=1$ over steps:
$\bm{x}_{t+\Delta t}=x_{t}+\Delta tf(\bm{x}_{t},t|\bm{o}),\qquad\text{for}\ t\in[0,1]$ (4)
3.2.3 Intriguing Properties of Flow Matching Policy
In this section, we provide some insights and intuitions about flow matching policy and its comparisons against diffusion policy for clarification. Algorithm 1 and Algorithm 2 respectively shows the pseudocode of training flow matching and diffusion policy with Denoising Diffusion Probabilistic Models (DDPM). We could see several differences between these two methods:
  1. 1.
    By solving a Stochastic Differential Equation (SDE), DDPM generates a clean sample from Gaussian noise. Flow matching regresses onto a target flow vector field that generates a deterministic mapping from source to target data distributions by solving an ordinary differential equation (ODE).
  2. 2.
    DDPM reverses a diffusion process that adds noise to a clean sample until it becomes Gaussian noise. Flow matching drops all the Gaussian assumptions.
  3. 3.
    Flow matching also includes a particularly interesting family of probability paths: the vector field that corresponds to the principle of Optimal Transport (OT) with linear interpolant. Paths with linear interpolation are simpler than diffusion paths, forming straight line trajectories whereas diffusion paths result in curved paths. These properties seem to empirically translate to more stable training, faster generation, and better performance.
Algorithm 1 Robot Flow Matching Policy Training
observation $\bm{o}$, target robot actions $\bm{x}_{1}$, source random waypoints $p_{0}$
flow $\mathbf{v}_{\bm{\theta}}$